10.10.10.100/28 answer the following question depending on the given ip address

A. network address

B.first host address

C.last host address

D.broadcast address


1
Expert's answer
2021-12-30T12:04:25-0500

A.

mask is /28, so last octet is: 11110000, so there are "2^4=16" bits for network (value of network address jumps through 16)

then  network address:

10.10.10.96


B.

first host address:

10.10.10.97


C.

last host address:

10.10.10.96 = 10.10.10.(96+16-2)=10.10.10.110


D.

broadcast address (last address of the network):

10.10.10.96 = 10.10.10.(96+16-1)=10.10.10.111